OpenStudy (anonymous):

Simplify root(4) (b^4).

OpenStudy (anonymous):

4th root of b^4 is b b*b*b*b = b^4 4th root means what times itself 4 times is the number. b*b*b*b=b^4

geerky42 (geerky42):

Not quite.

OpenStudy (anonymous):

Or, do you mean: sqrt(4)*b^4?

OpenStudy (anonymous):